# Modeling fibrous tissue in vascular fluid-structure interaction: a morphology-based pipeline and biomechanical significance

### v1

@@ 2.1.2 Tissue modeling and mesh generation
Based on whether the wall can be identified, the existing imaging modality can be categorized into two groups. The first group involves the computed tomography angiography and magnetic resonance angiography, both of which are widely used nowadays in pathology evaluations. A major limit is their lack of tissue contrast, making it difficult to identify the vascular and surrounding organs. The second group encompasses intravascular ultrasound, optical coherence tomography [40, 83], and black-blood magnetic resonance imaging (MRI) [6, 46], etc. Although they may provide detailed information about the vessel wall and allow direct segmentation, there are critical shortcomings that limit their availability, such as a restricted field of imaging for large vessels, susceptibility to motion artifacts, etc. We do not assume that the angiographic data contains vessel wall information, and the proposed procedures enable the preparation of a vascular tissue model based on most existing imaging modalities.
@@
We propose the following procedures for constructing the surface representation of the vascular wall. On each plane for 2D segmentation, the arithmetic mean of all contour points defines the centroid of the lumen at this cross section. We then generate a set of new contour points for the exterior wall surface by demanding them to be collinear with their corresponding luminal contour points and the centroid. The distance from the new contour points to the centroid is scaled according to the thickness-to-radius ratio $\delta$. The value of $\delta$ can be chosen according to physiological facts [17]. Alternatively, its value can also be adjusted on each plane to reflect local pathological characteristics. By interpolating the newly generated contour points, we obtain the closed B-spline curve delineating the exterior wall surface (Figure 1 (d) and (e)). The exterior wall surface can be generated by lofting the newly generated B-spline curves (Figure 1 (f) and (g)), and the smoothed discrete surface mesh can be created subsequently, following the procedures outlined above. The previously obtained luminal surface is directly adopted as the interior wall surface for the tissue.
@@
Given the exterior and interior surfaces of the tissue, we may close the volume by creating planar surfaces at an inlet and the outlets, which completes the B-Reps of the luminal and wall volumes. We refer to the surface on the proximal end of the artery as the inlet and the rest planar surfaces as outlets for the multi-branched tubular geometry. In specific, the tissue wall volume is bounded by the interior wall surface (a.k.a. the luminal surface), the exterior wall surface, and planar annular surfaces at an inlet and outlets. The luminal volume is bounded by the interior wall surface and the planar surfaces at the inlet and outlets. Different types of mesh generation algorithms can be used to complete the volumetric mesh [5, 28, 66]. The following outlines the specific steps for the vascular FSI mesh generation.
@@
    To resolve the near-wall viscous layer, a boundary resolving mesh is generated by extruding inward from the surface mesh of the lumen using the advancing layer method [25] with a specified number of layers and layer growth ratio. In particular, it is allowed to modify the luminal surface mesh to accommodate the boundary layer mesh generation in this step. This is crucial for the robust BL mesh generation. The updated surface mesh will remain unchanged for the subsequent steps.
@@
    Enclose the luminal volume by generating planar surface meshes at an inlet and outlets.
@@
    Generate the volumetric mesh for the lumen, as shown in Figure 2 (c).
@@
    Enclose the tissue volume by generating the planar annular surface meshes at the inlet and outlets. It is worth noting that the annular mesh may not be strictly coplanar with the corresponding planar surface mesh previously generated for the lumen. Therefore, one may need to edit the coordinates of the mesh nodes when necessary.
@@
    Generate the volumetric mesh for the vascular tissue, as depicted in Figure 2 (h), and ensure that the generated mesh strictly respect its B-Rep.
The two meshes together constitute a complete mesh for FSI analysis (Figure 2 (j)). We emphasize that mesh modification is only allowed in the first step, and the updated mesh is inherited for the subsequent steps. Critically, the generation of the tissue mesh does not allow the modification of the surface representation, which is crucial for ensuring the $C^{0}$-continuity of the mesh across the two domains, a necessary condition for proper coupling in FSI analysis.
@@
In contrast to traditional mesh extrusion techniques [51, 60, 80], the above approach exhibits greater robustness and precision. It generates the CAD model of the tissue surface by manipulating the contours of each 2D segmentation, followed by mesh generation, rather than directly manipulating lumen surface meshes to create tissue meshes. This effectively prevents self-intersection for complex geometric shapes with multiple branches. Furthermore, the above strategy allows convenient adjustment of local wall thickness.
@@
The idea of segmenting interior and exterior wall surfaces has been utilized for studying the correlation between hemodynamics and plaque progression based on black-blood MRI [46, 68]. Similar modeling approaches have been combined with FSI simulations for investigating atherosclerotic plaques in coronary and carotid arteries, using mutli-contrast MR [72] or intravascular ultrasound images [82]. Nevertheless, their emphasis was placed on the construction of the wall model with the plaques, and the geometries were relatively simple.
@@ 2.2 Definition of local basis vectors
To facilitate the generation of local basis vectors, we need to extract the centerline of the vascular tissue. For a vascular network with $N_{\mathrm{b}}$ branches, there exists $N_{\mathrm{b}}$ centerlines that extends from the inlet to an outlet. Solving the Eikonal equation yields the centerline curve $\bm{c}_{k}(s)$ and the radius $r_{k}(s)$ of the maximum inscribed sphere along this curve [58]. Here, $s\in[0,1]$ denotes the curvilinear abscissa along the centerline curve, and $1\leq k\leq N_{\mathrm{b}}$ is the branch index. Based on the above discussion, a scalar function $\theta_{k}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ associated with each centerline can be defined as
@@
Remark 3.
The centerline may also facilitate the generation of BL mesh with variable layer thicknesses. In specific, the radius information recorded along the centerline makes it convenient to adjust BL thicknesses to accommodate lumen diameter variations.
@@
Figure 4: The generation of local radial (a), pseudo-axial (b), and axial (c) vectors. (e) A multi-branched vessel. The red lines represent the centerlines.
@@
First, for a mesh node on the interior and exterior wall surfaces, the local radial direction can be defined by the normal vector at that node, pointing towards the exterior side of the wall. This can be achieved by algorithms developed for polygon surfaces [13, 43] or by solving a differential equation on the surface [18]. For a generic point $\bm{x}^{\mathrm{vol}}$ located inside of the vessel wall, we first traverse all mesh nodes on the interior and exterior wall surfaces to identify the closest nodes to $\bm{x}^{\mathrm{vol}}$, which are denoted by $\bm{x}^{\mathrm{in}}$ and $\bm{x}^{\mathrm{out}}$, respectively. The distance between $\bm{x}^{\mathrm{vol}}$ and $\bm{x}^{\mathrm{in}}$ ($\bm{x}^{\mathrm{out}}$) is denoted by $d^{\mathrm{in}}$ ($d^{\mathrm{out}}$). The local radial direction $\bm{e}_{\mathrm{r}}(\bm{x}^{\mathrm{vol}})$ at the point $\bm{x}^{\mathrm{vol}}$ is then defined as a weighted sum as
@@
When dealing with a single vessel, the previously mentioned strategy is adequate. Yet, for a vascular network with $N_{\mathrm{b}}$ centerlines, a point with the minimum distance can be found on each centerline. There is thus a need to define a unique associated point from these $N_{\mathrm{b}}$ points. A straightforward method to determine this is based on the absolute distance, that is, to identify the point with the closest distance to point $\bm{x}^{\mathrm{vol}}$ among these $N_{\mathrm{b}}$ points as the associated point. Nevertheless, when two branches are close to each other, such as the branches with centerlines $\bm{c}_{1}^{h}(s)$ and $\bm{c}_{2}^{h}(s)$ depicted in Figure 5, the closest point $\bm{x}^{1}$ on the centerline $\bm{c}^{h}_{1}(s)$ is identified as the associated point when $d^{1}<d^{2}$, which results in the generation of non-physiological local basis vectors. To address this issue, we propose a refined metric that utilizes the local radius information recorded on the centerline points. Specifically, for each of the $N_{\mathrm{b}}$ points, the relative distance ${d^{k}}/{r^{k}}$ to the point $\bm{x}^{\mathrm{vol}}$ is calculated. The value of $r^{k}$ is the radius at the closest point on the $k$-th centerline, and it is at a distance of $d^{k}$ from the point $\bm{x}^{\mathrm{vol}}$. Typically, the relative distance is significantly greater than $1.0$ for points on unrelated centerlines. Consequently, the point with the smallest relative distance is recognized as the associated point. The detailed implementation of this procedure is presented in Algorithm 1.
@@
3:All discrete centerlines $\{\bm{c}^{h}_{k}(s)\}_{k=1}^{N_{\mathrm{b}}}$ with their centerline point sets $\{\bm{x}_{i}^{k}\}_{i=1}^{N_{k}}$ and radius sets $\{r_{i}^{k}\}_{i=1}^{N_{k}}$
@@
The inherent challenge in generating local basis vectors for multi-branched vessels based on centerlines lies in identifying the correct centerline. In addition to the proposed relative distance method, an alternative approach was presented in [62], in which a cone defined by a prescribed angle was utilized to select the correct centerline.
@@
In most cases when dealing with multi-branched vessels, the application of the aforementioned relative distance proves effective. However, challenges arise near bifurcation regions when one branch is significantly larger than the other. This may leads to a tendency of misidentifying the associated point. We elaborate on this situation and propose a refined strategy to address it in the following text.
@@
The bifurcation region is defined as the location where one centerline intersects the canal surface of another centerline. For the two centerlines $\bm{c}^{h}_{1}(s)$ and $\bm{c}^{h}_{2}(s)$ that constitute a bifurcation region, we may identify one as the parent and the other one as the child branch according to their local radii near the bifurcation. Let the intersection point between the canal surface of $\bm{c}^{h}_{1}(s)$ and the centerline $\bm{c}^{h}_{2}(s)$ be $\bm{x}_{1}$. Along the centerline $\bm{c}^{h}_{1}(s)$, there exists a sphere that touches the point $\bm{x}_{1}$ due to the definition of the canal surface, and we denote the radius of the sphere as $r_{1}$. Conversely, the intersection point between the canal surface of $\bm{c}^{h}_{2}(s)$ and the centerline $\bm{c}^{h}_{1}(s)$ is denoted by $\bm{x}_{2}$, and the radius of the corresponding sphere is denoted as $r_{2}$. With these local geometric information, we may identify $\bm{c}^{h}_{1}(s)$ and $\bm{c}^{h}_{2}(s)$ as the parent and child branches if $r_{1}>r_{2}$; otherwise, if $r_{1}<r_{2}$, $\bm{c}^{h}_{1}(s)$ is the child branch, and $\bm{c}^{h}_{2}(s)$ is the parent branch.
@@
When invoking the Eikonal Equation, the local radius of the child branch smoothly transits from large to small values near the bifurcation region. This causes a non-physiological overestimation of the local radius, as shown in Figure 6 (a). This issue is visualized by inspecting the canal surface of different centerlines shown in Figure 6 (b). As a result, there is a tendency to locate the associated point on the child branch near the bifurcation when using the relative distance, causing incorrect local basis vector assignment on the parent branch, as illustrated in Figure 6 (c).
@@
To resolve this issue, we first identify if the child centerline point is located within the canal surface of the parent centerline. This can be achieved by evaluating the relative distance of the point to the parent centerline. When these points are identified (i.e. the relative distance is less than $1.0$), their recorded radius is reduced by a scaling function $\phi$, which is a univariate function of the relative distance. The function $\phi$ is a smooth, monotonically increasing function and satisfies $\phi(1)=1$ and $\phi(0)=\varepsilon$. Here, $\varepsilon$ is a small positive number, ensuring $\phi(0)$ is not zero to avoid division-by-zero errors in subsequent calculation of the relative distance. The specific criteria for identifying the bifurcation region and the algorithm for reducing the radii from the child branches are detailed in Algorithm 2. As shown in Figure 6 (d), the local radii recorded on the child branch at the bifurcation are reduced, and in Figure 6 (e) the refined tube surfaces also indicate that the erroneous influence of the child branch has been corrected. By employing the relative distance at the refined bifurcation region, the correct local basis vectors are generated (Figure 6 (f)). We mention that for a set of multiple centerlines, these procedures have to be performed for each bifurcation region.
Algorithm 2 Update radii of one child branch centerline points
@@
We observed that for vascular models with bifurcations, defining local basis vectors at mesh nodes inevitably results in the presence of two pairs of oppositely local circumferential and local axial basis vectors at the nodes of elements near the bifurcation (see the element ➀ in Figure 7). In such cases, if we calculate the local circumferential and axial basis vectors at the quadrature points in these elements through interpolation, the definition of the local basis vectors may appear singular, causing divergence in the calculation. It is thus recommended generating the local basis vectors at the quadrature points.

### v2

@@ 2.1.2 Tissue modeling and mesh generation
Based on whether the wall can be identified, the existing imaging modality can be categorized into two groups. The first group involves the computed tomography angiography and magnetic resonance angiography, both of which are widely used nowadays in pathology evaluations. A major limit is their lack of tissue contrast, making it difficult to identify the vascular and surrounding organs. The second group encompasses intravascular ultrasound, optical coherence tomography [40], and black-blood magnetic resonance imaging (MRI) [6], etc. Although they may provide detailed information about the vessel wall and allow direct segmentation, there are critical shortcomings that limit their availability, such as a restricted field of imaging for large vessels, susceptibility to motion artifacts, etc. We assume that the angiographic data does not contain vessel wall information, and the proposed procedures enable the preparation of a vascular tissue model based on most existing imaging modalities.
@@
We propose the following procedures for constructing the surface representation of the vascular wall. On each plane for 2D segmentation, the arithmetic mean of all contour points defines the centroid of the lumen at this cross-section. We then generate a set of new contour points for the exterior wall surface by demanding them to be collinear with their corresponding luminal contour points and the centroid. The distance from the new contour points to the centroid is scaled according to the thickness-to-radius ratio $\delta$. The value of $\delta$ can be chosen according to physiological facts [17, 41]. Alternatively, its value can also be adjusted on each plane to reflect local pathological characteristics. By interpolating the newly generated contour points, we obtain the closed B-spline curve delineating the exterior wall surface (Figure 1 (d) and (e)). The exterior wall surface can be generated by lofting the newly generated B-spline curves (Figure 1 (f)), and the smoothed discrete surface mesh can be created subsequently, following the procedures outlined above. The previously obtained luminal surface is directly adopted as the interior wall surface for the tissue.
@@
Given the exterior and interior surfaces of the tissue, we may close the volume by creating planar surfaces at the inlet and outlets, which completes the B-Reps of the luminal and wall volumes. Hereafter, we refer to the planar surface on the proximal end of the artery as the inlet and the rest planar surfaces as outlets for the multi-branched tubular geometry. In specific, the tissue wall volume is bounded by the interior wall surface (a.k.a. the luminal surface), the exterior wall surface, and planar annular surfaces at an inlet and outlets. The luminal volume is bounded by the interior wall surface and the planar surfaces at the inlet and outlets. Mesh generation algorithms can be used to complete the volumetric mesh [5, 28, 66]. The following outlines the specific steps for the vascular FSI mesh generation.
@@
    To resolve the near-wall viscous layer, a boundary resolving mesh is generated by extruding inward from the surface mesh of the lumen using the advancing layer method [25] with a specified number of layers and layer growth ratio. In particular, it is allowed to modify the luminal surface mesh to accommodate the boundary layer mesh generation in this step. This is crucial for robust BL mesh generation. The updated surface mesh will remain unchanged for the subsequent steps.
@@
    Enclose the luminal volume by generating planar surface meshes at the inlet and outlets.
@@
    Generate the volumetric mesh for the lumen (Figure 2 (a)).
@@
    Enclose the tissue volume by generating the planar annular surface meshes at the inlet and outlets.
@@
    Generate the volumetric mesh for the vascular tissue (Figure 2 (b)), and ensure that the generated mesh strictly respects its B-Rep.
The two meshes together constitute a complete mesh for FSI analysis (Figure 2 (c)). We emphasize that mesh modification is only allowed in the first step, and the updated mesh is maintained for the subsequent steps. This is crucial for ensuring the $C^{0}$-continuity of the mesh across the two domains, a necessary condition for proper coupling in FSI analysis.
Figure 2: (a) Generating the volumetric mesh for the lumen; (b) closing the annular planar surfaces to form the vessel wall, followed by the generation of its volumetric mesh; (c) combining the meshes to obtain the vascular FSI mesh.
@@
In contrast to mesh extrusion techniques [51, 60, 79], the above approach exhibits greater robustness and precision. It generates the CAD model of the tissue surface by manipulating the contours of each 2D segmentation, followed by mesh generation, rather than directly manipulating lumen surface meshes to create tissue meshes. This effectively prevents self-intersection for complex geometric shapes with multiple branches. Furthermore, the above strategy allows convenient adjustment of local wall thickness.
@@
The idea of segmenting interior and exterior wall surfaces has been utilized for studying the correlation between hemodynamics and plaque progression based on black-blood MRI [46, 68]. Similar modeling approaches have been combined with FSI simulations for investigating atherosclerotic plaques in coronary and carotid arteries, using multi-contrast MR [72] or intravascular ultrasound images [81]. Nevertheless, their emphasis was placed on the construction of the wall model with the plaques, and the geometries were relatively simple.
@@ 2.2 Definition of local basis vectors
To facilitate the generation of local basis vectors, we need to extract the centerline of the vascular tissue. For a vascular network with $N_{\mathrm{b}}$ branches, there exists $N_{\mathrm{b}}$ centerlines that extend from the inlet to an outlet. Solving the Eikonal equation yields the centerline curve $\bm{c}_{k}(s)$ and the radius $r_{k}(s)$ of the maximum inscribed sphere within the exterior wall along this curve [58]. Here, $s\in[0,1]$ denotes the curvilinear abscissa along the centerline curve, and $1\leq k\leq N_{\mathrm{b}}$ is the branch index. Based on the above, a scalar function $\theta_{k}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ associated with each centerline can be defined as
@@
Figure 4: The generation of local (a) & (b) radial, (c) pseudo-axial, and (d) axial vectors. (e) The red lines depict the centerlines of a multi-branched vascular model.
@@
First, for a mesh node on the interior or exterior wall surfaces, the local radial direction can be defined by the normal vector at that node, pointing towards the exterior side of the wall. This is achieved by algorithms developed for polygon surfaces [13, 43] or by solving a differential equation on the surface [18]. For a generic point $\bm{x}^{\mathrm{vol}}$ located inside of the vessel wall, we first traverse all mesh nodes on the interior and exterior wall surfaces to identify the closest nodes to $\bm{x}^{\mathrm{vol}}$, which are denoted by $\bm{x}^{\mathrm{in}}$ and $\bm{x}^{\mathrm{out}}$, respectively. The distance between $\bm{x}^{\mathrm{vol}}$ and $\bm{x}^{\mathrm{in}}$ ($\bm{x}^{\mathrm{out}}$) is denoted by $d^{\mathrm{in}}$ ($d^{\mathrm{out}}$). The local radial direction at the point $\bm{x}^{\mathrm{vol}}$ is then defined as a weighted sum as
@@
When dealing with a single vessel, the previously mentioned strategy is adequate. Yet, for a vascular network with $N_{\mathrm{b}}$ centerlines, a point with the minimum distance can be found on each centerline. There is thus a need to define a unique associated point from these $N_{\mathrm{b}}$ points. A straightforward method is based on the absolute distance, that is, to identify the point with the closest distance to point $\bm{x}^{\mathrm{vol}}$ among these $N_{\mathrm{b}}$ points as the associated point. Nevertheless, when two branches are close to each other, such as the branches with centerlines $\bm{c}_{1}^{h}(s)$ and $\bm{c}_{2}^{h}(s)$ depicted in Figure 5, the closest point $\bm{x}^{1}$ on the centerline $\bm{c}^{h}_{1}(s)$ is identified as the associated point when $d^{1}<d^{2}$, which results in the generation of non-physiological local basis vectors. To address this issue, we propose a refined metric that utilizes the local radius information recorded on the centerline points. Specifically, for each of the $N_{\mathrm{b}}$ points, the relative distance ${d^{k}}/{r^{k}}$ to the point $\bm{x}^{\mathrm{vol}}$ is calculated. The value of $r^{k}$ is the radius at the closest point on the $k$-th centerline, and it is at a distance of $d^{k}$ from the point $\bm{x}^{\mathrm{vol}}$. The relative distance will be significantly greater than $1.0$ for points on unrelated centerlines. Consequently, the point with the smallest relative distance is recognized as the associated point. The implementation of this procedure is detailed in Algorithm 1.
@@
3:All discrete centerlines with their centerline point sets $\{\bm{x}_{i}^{k}\}_{i=1}^{N_{k}}$ and radius sets $\{r_{i}^{k}\}_{i=1}^{N_{k}}$
@@
The inherent challenge in generating local basis vectors for multi-branched vessels based on centerlines lies in identifying the correct centerline. An alternative approach was presented in [62], in which a cone defined by a prescribed angle was utilized to select the correct centerline.
@@
Near bifurcation regions when one branch is significantly larger than the other, using the relative distance leads to a tendency to misidentify the associated point. We elaborate on this situation and propose a refined strategy to address it in the following.
@@
The bifurcation region is defined as the location where one centerline intersects the canal surface of another centerline. Consider two centerlines $\bm{c}^{h}_{1}(s)$ and $\bm{c}^{h}_{2}(s)$ that constitute a bifurcation region. We may identify one as the parent and the other one as the child branch according to their local radii near the bifurcation. Let the intersection point between the canal surface of $\bm{c}^{h}_{1}(s)$ and the centerline $\bm{c}^{h}_{2}(s)$ be $\bm{x}_{1}$. Along the centerline $\bm{c}^{h}_{1}(s)$, there exists a sphere that touches the point $\bm{x}_{1}$ due to the definition of the canal surface, and we denote the radius of the sphere as $r_{1}$. Conversely, the intersection point between the canal surface of $\bm{c}^{h}_{2}(s)$ and the centerline $\bm{c}^{h}_{1}(s)$ is denoted by $\bm{x}_{2}$, and the radius of the corresponding sphere is denoted as $r_{2}$. With these local geometric information, we may identify $\bm{c}^{h}_{1}(s)$ and $\bm{c}^{h}_{2}(s)$ as the parent and child branches if $r_{1}>r_{2}$; otherwise, if $r_{1}<r_{2}$, $\bm{c}^{h}_{1}(s)$ is the child branch, and $\bm{c}^{h}_{2}(s)$ is the parent branch.
@@
Due to the nature of the Eikonal Equation, the local radius of the child branch smoothly transits from large to small values near the bifurcation region. This causes a non-physiological overestimation of the local radius, as shown in Figure 6 (a). This issue is visualized by inspecting the canal surface of different centerlines shown in Figure 6 (b). As a result, there is a tendency to locate the associated point on the child branch near the bifurcation when using the relative distance, causing incorrect local basis vector assignment on the parent branch, as illustrated in Figure 6 (c).
@@
To resolve this issue, we first identify if the child centerline point is located within the canal surface of the parent centerline. This can be achieved by evaluating the relative distance of the point to the parent centerline. When these points are identified (i.e. points with the relative distance less than $1.0$), their recorded radius is reduced by a scaling function $\phi$, which is a univariate function of the relative distance. The function $\phi$ is a smooth, monotonically increasing function and satisfies $\phi(1)=1$ and $\phi(0)=\varepsilon$. Here, $\varepsilon$ is a small positive number, ensuring $\phi(0)$ is not zero to avoid division-by-zero errors in the subsequent calculation of the relative distance. The specific criteria for identifying the bifurcation region and the algorithm for reducing the radii from the child branches are detailed in Algorithm 2. As shown in Figure 6 (d), the local radii recorded on the child branch at the bifurcation are reduced, and the refined tube surfaces also indicate that the erroneous influence of the child branch has been corrected (Figure 6 (e)). Employing the relative distance with the radius values modified, the correct local basis vectors are generated (Figure 6 (f)). We mention that for a set of multiple centerlines, these procedures have to be performed for each bifurcation region.
Algorithm 2 Modification of the radii of one child branch centerline points
@@
It is possible to have two pairs of local basis vectors pointing in opposite directions at the mesh nodes of an element near the bifurcation region (e.g., see the element ➀ in Figure 7). In such a case, if we calculate the local circumferential and axial basis vectors at the quadrature points through interpolation, the definition of the local basis vectors will become singular. In our practice, the local basis vectors are generated at the quadrature points directly.